# 3M-Health: Multimodal Multi-Teacher Knowledge Distillation for Mental Health Detection

### v2

@@ 5.5. Impact of Student Model Inputs
Table 10. Ablation study using different combinations of input modalities to the student model. Bold face indicates best score while second best are underlined. A ✓indicates the addition of the emotion-based (Emo) and/or the audio-based (Aud) input features. Highlighted rows show our original proposed student setup. VT: randomly initialised vanilla transformer.
@@
We examine different combinations of multimodal inputs for the student model in Table 10 in order to explore the optimal input for a knowledge distillation for the student model. In this scenario, we concatenate emotion embeddings, audio embeddings, or both with textual post embeddings from pre-trained BERT and then pass them to the transformer layer after a linear layer projection. We also test a randomly initialised vanilla transformer compared to pre-trained BERT for the student model. The results indicate that unimodal textual post inputs outperform the concatenation of multimodal inputs for the student model. Moreover, pre-trained BERT yields better results than the randomly initialised vanilla transformer across all datasets. These outcomes underscore the effectiveness of multi-aspect knowledge acquired from the multi-teachers, efficiently guiding the student to achieve robust performance with only textual inputs.
5.6. Hyperparameter Testing
We further investigate the impact of different patch size values for the audio-based teacher model, Audio Spectrogram Transformer (AST) (Gong et al., 2021), while maintaining a consistent setup for the student model. Figure 6 shows each dataset’s weighted F1 score for each patch size value tested. TwitSuicide, DEPTWEET, and IdenDep datasets show a relatively stable performance between 2 to 64 patch sizes; however, for SDCNL, performance improvement may be achieved using a patch size of 32. This improvement may be due to the higher variance in audio duration of the outliers in the SDCNL dataset compared to the other three datasets, as shown in Figure 3. Despite being shorter on average length and duration than the other Reddit-based dataset, SDCNL has audio samples with longer durations, which may have benefited from a patch size 32. However, a sharp decline in performance could be expected when the patch size is increased to 64.

### v3

@@ 5.5. Impact of Student Model Inputs
Table 10. Ablation study using different combinations of input modalities to the student model. Bold face indicates best score while second best are underlined. A ✓indicates the addition of the emotion-based (Emo) and/or the audio-based (Aud) input features. Highlighted rows show our proposed student setup. VT: randomly initialised vanilla transformer.
@@
We examine different combinations of multimodal inputs for the student model in Table 10 in order to explore the optimal input for a knowledge distillation for the student model. We concatenate emotion embeddings, audio embeddings, or both with text embeddings from pre-trained BERT and then pass them to the transformer layer after a linear layer projection. We also test a randomly initialised vanilla transformer compared to the pre-trained BERT. The results indicate that unimodal textual post inputs outperform the concatenation of multimodal inputs for the student model. Moreover, pre-trained BERT yields better results than the randomly initialised vanilla transformer across all datasets. These outcomes underscore the effectiveness of multi-aspect knowledge acquired from the multi-teachers, efficiently guiding the student to achieve robust performance with only textual inputs.
5.6. Hyperparameter Testing
We further investigate the different patch size values for the audio-based teacher model while maintaining a consistent setup for the student model. Figure 6 shows each dataset’s weighted F1 score for each patch size value. TwitSuicide, DEPTWEET, and IdenDep datasets show a relatively stable performance between 2 to 64 patch sizes; however, for SDCNL, performance improvement may be achieved using a patch size of 32. This may be due to the higher variance in audio duration of the outliers compared to the other three datasets (Figure 3). Despite being shorter on average length and duration than the other Reddit-based dataset, SDCNL has some longer audio samples, which may have benefited from a patch size 32. However, a sharp decline in performance could be expected when the patch size is increased to 64.
@@ 6. Conclusion
Acknowledgements.
This study was supported by funding from the Google Award for Inclusion Research Program (G222897).